OpenStudy (wwe123):

2+7tan²ø = 3.25sec²ø for 0<=ø>=2p

OpenStudy (anonymous):

use the identity\[\sec^2 \theta=1+\tan ^2 \theta\]

OpenStudy (anonymous):

can u solve it from here?

OpenStudy (anonymous):

\[2+7\tan^2 \theta = 3.25(1+\tan^2 \theta)\]\[\tan^2 \theta =\frac{1}{3}\]